OpenStudy (anonymous):

Solve the giving quadratic equation by factoring: x/2 + 1/x-3 = 3

OpenStudy (anonymous):

I'm helping my friend on this, and we keep getting to this point: \[(x ^{2} - 9x + 20)\div (2(x-3))\] and then we get lost. I've gone over it a few times to make sure that the step we're on is correct, but we don't know where to go from here.

OpenStudy (dumbcow):

not sure where you get to that point, let me work out the steps and then see where you go wrong \[\frac{x(x-3)}{2(x-3)}+\frac{2}{2(x-3)} = 3\] \[\frac{x^{2}-3x+2}{2(x-3)} = 3\] \[x^{2}-3x+2=6x-18\] \[x^{2}-9x+20 = 0\] \[(x-5)(x-4) = 0\]

OpenStudy (anonymous):

Oh man, we started it off the right way, but got confused and convinced ourselves that we had to bring the 3 over at the start. Thanks for the help, another job well done dumbcow, you're the bomb!

OpenStudy (dumbcow):

thanks oh i see, what you had above equaled zero then that would still work, just set the numerator equal to 0 and you get same answer :)